Note:

1 mole of a compound is equivalent to its molecular mass

1 mole of a compound is equal to 6.02x10^23 molecules

The percentage is a direct equivalent to grams

Answers should be in 3 significant figures

No need to include the units and the compound in the answers


  1. Determine the molecular formula of a compound with an empirical formula of NH2 and a formula mass of 32.06 u. 
  2. A sample of a compound with a formula mass of 34.00 u is found to consist of 0.44 g H and 6.92 g O. Find its molecular formula. 
1
Expert's answer
2021-06-28T08:11:13-0400

1.

NH2 have emperical mass .


14 + 2 = 16 .



Given molecular mass = 32 .


n = 32 / 16 = 2 .


Molecular formula = 2 ( NH2) = N2H4 .



2.


Given

moles of H = 0.44 / 1 = 0.44 mole .

moles of O = 6.92 / 16 = 0.44 mole .

molar ratio of H and O = 1: 1

So, emperical formula = HO .

Emperical mass = 1 + 16 = 17 .

Given molar mass = 34 .

n = 34 / 17 = 2 .

Molecular formula = 2 ( HO )

= H2O2 .